### v2

# Flow between Extremal One-Point Energy Correlators in QCD

Abstract
The energy density generated by a vector current is characterized by a single parameter $a_{\mathcal{E}}$ bounded by unitarity to $1/2\leq a_{\mathcal{E}}\leq 1$, with extremal values saturated by free theories of different matter content. Through confinement, QCD transmutes fermionic matter into scalars, revealing a nontrivial flow between extremal correlators. We reconstruct this flow using perturbative QCD and chiral perturbation theory. The observable is accessible with currently available experimental data.
preprint: CERN-TH-2025-175
On its 50th anniversary Gross and others (2023), quantum chromodynamics (QCD) remains a vibrant and rich field, a whetstone for theoretical physicists to sharpen their understanding of quantum field theory and of nature. A defining feature of QCD is confinement, which dramatically changes dynamics as it flows from the ultraviolet (UV) to the infrared (IR). Observables appropriate in one regime often lose their utility in the other: the $S$ matrix efficiently characterizes two-pion scattering, but quickly becomes unfeasible once multiparticle thresholds open, while jet rates are well defined at high energies but not at low.
Event shapes provide a class of observables that can be studied across scales. Indeed, it was through their study that a first understanding of gluon emission emerged Ellis et al. (1976); Farhi (1977); Georgi and Machacek (1977); Fox and Wolfram (1978). A particularly relevant set of event shapes are energy correlators Basham et al. (1978a, b), which possess rich theoretical and phenomenological properties, recently reviewed in Moult and Zhu (2025). While much of the phenomenological studies have been devoted to multipoint correlators, the simplest case, the one-point energy correlator, remains largely unexplored.
The angular dependence of the one-point energy correlator of a vector current is fully characterized by a single parameter $a_{\mathcal{E}}$
$\langle{\mathcal{E}}_{n}\rangle\,=\,\frac{\langle{\mathcal{E}}\rangle}{4\pi}\left[1+a_{\mathcal{E}}\left(\frac{3}{2}\sin^{2}\theta-1\right)\right]~,$ (1)
with $\langle{\mathcal{E}}\rangle$ the total energy. The parameter $a_{\mathcal{E}}$ is constrained by energy positivity to lie within $1/2\leq a_{\mathcal{E}}\leq 1$ Hofman and Maldacena (2008). While free theories with different matter content saturate the different extremal values of $a_{\mathcal{E}}$ Zhiboedov (2014), strongly coupled theories that transmute their degrees of freedom may transition from one boundary to the other. In this Letter, we observe that QCD provides a striking realization of this phenomenon, flowing between the two extremal energy correlators as the theory converts fermionic degrees of freedom in the UV into bosonic ones in the IR.
The goal of this Letter is to determine the global evolution of $a_{\mathcal{E}}$ from the electroweak scale all the way down to the pion threshold by combining the predictions from perturbative QCD with the ones of chiral perturbation theory. Our main result is summarized in Figure 1. The remainder of the Letter is devoted to outlining the theoretical framework and key ingredients that lead to it.
Flowing one-point energy correlator.   The one-point energy correlator measured along $\vec{n}$ on a state sourced by the electromagnetic vector current $J^{\mu}$ is defined from the three-point correlator between the currents and the energy flow operator $\mathcal{E}_{n}$
$H^{\mu\nu}_{\mathcal{E}}\,=\,\int d^{4}xe^{iq\cdot x}\langle J^{\mu\dagger}(x)\mathcal{E}_{n}J^{\nu}(0)\rangle~.$ (2)
Dotting it to a polarization basis along the current spin axis allows to define the density matrix $\langle{\mathcal{E}}_{n}\rangle_{h^{\prime}h}\,=\,\epsilon^{*\mu}_{h^{\prime}}H^{\mu\nu}_{\mathcal{E}}\epsilon^{\nu}_{h}$. Given that the correlator only depends on the total injected momentum $q^{\mu}$ and the detector direction $n^{\mu}$, in the rest frame with $q^{\mu}=(\sqrt{q^{2}},\vec{0})$ and $n^{\mu}=(1,\vec{n})$, the density matrix is fully determined in terms of a single parameter $a_{\mathcal{E}}$
$\langle{\mathcal{E}}_{n}\rangle_{h^{\prime}h}=\frac{\langle{\mathcal{E}}\rangle}{4\pi}\left[\delta_{h^{\prime}h}+a_{\mathcal{E}}\left(3(n\cdot\epsilon^{*}_{h^{\prime}})(n\cdot\epsilon_{h})-\delta_{h^{\prime}h}\right)\right]~,$ (3)
where $\langle{\mathcal{E}}\rangle=\sqrt{q^{2}}$ is the total energy injected in the current and $a_{\mathcal{E}}$ measures the relative weight between the inclusive and the spinning part of the correlator.
For the unpolarized case of $e^{+}e^{-}\to\gamma^{*}\to\text{hadrons}$, the beam produces the trace of transverse polarizations $\frac{1}{2}{\rm tr\,}\langle{\mathcal{E}}_{n}\rangle_{h^{\prime}h}$, which implies $\delta_{h^{\prime}h}\to 1$ and $(n\cdot\epsilon^{*}_{h^{\prime}})(n\cdot\epsilon_{h})\to\frac{1}{2}\sin^{2}\theta$ in Eq. 3, where $\theta$ is the angle between spin axis of the current, in this case coinciding with the beam axis, and the detector direction. The explicit form of the one-point energy correlator is given in Eq. 1.
The two independent Lorentz scalars that can be built out of the three point correlator in Eq. 2 are $\eta_{\mu\nu}H^{\mu\nu}_{\mathcal{E}}$ and $n_{\mu}n_{\nu}H^{\mu\nu}_{\mathcal{E}}$. They determine $a_{\mathcal{E}}$ as
$a_{\mathcal{E}}\,=\,-\frac{1}{2}\left(1-3\,\frac{n_{\mu}n_{\nu}H^{\mu\nu}_{\mathcal{E}}}{-\eta_{\mu\nu}H^{\mu\nu}_{\mathcal{E}}}\right)~,$ (4)
depending on the total energy $q^{2}$. The unitarity and positivity of the energy imply that the contractions of the correlator obey the constraints $n_{\mu}n_{\nu}H^{\mu\nu}_{\mathcal{E}}\geq 0$ and $\eta_{\mu\nu}H^{\mu\nu}_{\mathcal{E}}-n_{\mu}n_{\nu}H^{\mu\nu}_{\mathcal{E}}\geq 0$, which imply the bounds $1/2\,\leq\,a_{\mathcal{E}}\,\leq\,1\,$, for the $a_{\mathcal{E}}$ parameter controlling the shape of the energy density Hofman and Maldacena (2008). The $a_{\mathcal{E}}$ depends on the average angular momentum of the generated state along the $\vec{n}$ direction. Upper and lower bounds are given by states with $m=0$ and $|m|=1$ spin along $\vec{n}$, respectively. Examples include a current of minimally coupled fermions ($a_{\mathcal{E}}=-1/2$), scalars ($a_{\mathcal{E}}=1$), or a Wess-Zumino-Witten (WZW) term $\epsilon^{\mu\nu\rho\sigma}\partial_{\nu}\pi^{a}\partial_{\rho}\pi^{b}\partial_{\sigma}\pi^{c}$ ($a_{\mathcal{E}}=-1/2$).
In the following, we present the calculation of the evolution of $a_{\mathcal{E}}$ as a function of the momentum $q^{2}$, tracking its flow from the UV to the IR.
—Perturbative QCD.  We can compute the correlator in Eq. 2 in perturbation theory by noting that the energy operator ${\mathcal{E}}_{n}$ acts on multiparticle states $|\alpha\rangle$ as ${\mathcal{E}}_{n}|\alpha\rangle=\sum_{i}E_{i}\delta^{(2)}(\Omega_{n}-\Omega_{i})|\alpha\rangle$. This leads a perturbative representation of the correlator as squared matrix elements with a phase space weighted by the energy and a sum over all states with particles in the direction $\vec{n}$.
The perturbative determinations of $\eta_{\mu\nu}H^{\mu\nu}_{\mathcal{E}}$ and $n_{\mu}n_{\nu}H^{\mu\nu}_{\mathcal{E}}$ can actually be derived from the computations of the longitudinal and transverse fragmentation functions in $e^{+}e^{-}$ annihilation Nason and Webber (1994). The semi-inclusive process $e^{+}e^{-}\to hX$, where $h$ represents a given hadron species and $X$ is the rest of the state, is described by separating the cross section into a transverse ($T$), longitudinal ($L$), and asymmetric ($A$) part, each one having an inclusive cross section of the form
$\frac{d\sigma^{h}_{P}}{dx}=\sum_{i}\int_{x}^{1}\frac{dz}{z}C_{P,i}(z,q^{2},\mu)D_{i}^{h}(x/z,\mu)~,$ (5)
where P=T,L,AP=T,\,L,\,A and $i$ runs over the types of partons. $C_{P,i}$ is calculable in perturbation theory and $D_{i}^{h}$ is the so-called fragmentation function for $h$ from $i$ partons. In the energy correlator, one is fully inclusive over $h$ and the type of parton $i$, which implies sensitivity only to the hard coefficients $C_{P,i}$. Indeed, using the fact that $\sum_{h}\int_{0}^{1}dzzD^{h}_{i}(z,\mu)=1$ due to the energy conservation, this allows us to represent tensor structures of the energy correlator at high energies in terms of integrals of $C_{P,i}$,
$-\eta_{\mu\nu}H^{\mu\nu}_{\mathcal{E}}$ $=$ $\frac{1}{2}\sum_{i}\int_{0}^{1}dx\,x\left[C_{T,i}(x,q^{2})+C_{L,i}(x,q^{2})\right]~,$
$n_{\mu}n_{\nu}H^{\mu\nu}_{\mathcal{E}}$ $=$ $\frac{1}{2}\sum_{i}\int_{0}^{1}dx\,x\,C_{L,i}(x,q^{2})~.$ (6)
The factor $x$ in the integrand corresponds to the energy weight, and it annihilates the soft part of the coefficient functions. The collinear divergences given by $1/\epsilon$ terms in $C_{P,i}$ are proportional to the parton splitting functions. They vanish after being inclusive over the partons $i$, implied by the collinear safety of the observable. The asymmetric part does not enter in the one-point energy correlator, but it does enter in the parity-odd terms of the one-point charge correlator of an electroweak current Riembau and Son (2025).
It is interesting to realize that at Born level $n_{\mu}n_{\nu}H^{\mu\nu}_{\mathcal{E}}=0$, and therefore, the boundary $a_{\mathcal{E}}=-1/2$ is saturated. This is equivalent to the Callan-Gross relation, which implies that at high energies the electromagnetic current is made out of only spin-1/2 fields Callan and Gross (1969). At next-to-leading order (NLO), the correction is the same for both tensor structures, given by $\frac{3}{2}(\alpha_{s}/2\pi)C_{F}\,$ Basham et al. (1978a, b). At NNLO, the corrections for $C_{P,i}$ were obtained in Rijken and van Neerven (1997, 1996); Mitov and Moch (2006); Blumlein and Ravindran (2006). The $\alpha_{s}^{3}$ calculation of the hard coefficient functions $C_{P,i}$ has been recently achieved in He et al. (2025). By taking the latter results and performing the integrals, with the help of the packages HPL and PLT Maitre (2006); Duhr and Dulat (2019), we get the $\alpha_{s}^{3}$ expressions for $\eta_{\mu\nu}H^{\mu\nu}_{\mathcal{E}}$ and $n_{\mu}n_{\nu}H^{\mu\nu}_{\mathcal{E}}$ reported in the Appendix. As a validation, the expression for $\eta_{\mu\nu}H^{\mu\nu}_{\mathcal{E}}$ matches the numerical expression for the $\mathcal{O}(\alpha_{s}^{3})$ term for the hadronic $R$ ratio in Baikov et al. (2012). Expanding the two Lorentz contractions in Eq. 4 up to order $\alpha_{s}^{3}$, it leads to the N3LO expression for the parameter $a_{\mathcal{E}}$ in perturbative QCD for massless quarks
$\begin{split}a_{\mathcal{E}}\,\,=\,\,&-\frac{1}{2}\,+\,\frac{\alpha_{s}}{\pi}\frac{9C_{F}}{8}\,+\,\left(\frac{\alpha_{s}}{\pi}\right)^{2}\left[-C_{F}^{2}\frac{99}{64}+C_{F}C_{A}\left(\frac{2023}{320}-\frac{9}{20}\zeta_{3}\right)-C_{F}n_{F}\frac{37}{32}\right]\\ &+\left(\frac{\alpha_{s}}{\pi}\right)^{3}\bigg[\,\,\,C_{F}^{3}\left(-\frac{3}{1280}-\frac{537}{40}\zeta_{3}+\frac{69}{4}\zeta_{5}\right)+C_{F}^{2}C_{A}\left(-\frac{14737}{960}-\frac{165}{16}\zeta_{2}+\frac{3803}{160}\zeta_{3}+\frac{605}{32}\zeta_{4}-\frac{231}{8}\zeta_{5}\right)\\ &\hskip 51.21504pt+C_{F}C_{A}^{2}\left(\frac{2599751}{57600}+\frac{319}{64}\zeta_{2}-\frac{307}{25}\zeta_{3}-\frac{209}{20}\zeta_{4}+\frac{87}{16}\zeta_{5}\right)\\ &\hskip 51.21504pt+C_{F}^{2}n_{f}\left(\frac{6163}{3840}+\frac{15}{8}\zeta_{2}-\frac{71}{40}\zeta_{3}-\frac{55}{16}\zeta_{4}+3\zeta_{5}\right)+C_{F}n_{f}^{2}\left(\frac{461}{360}-\frac{1}{16}\zeta_{2}+\frac{3}{40}\zeta_{3}\right)\\ &\hskip 51.21504pt+C_{F}C_{A}n_{f}\left(-\frac{3622}{225}-\frac{9}{16}\zeta_{2}+\frac{259}{400}\zeta_{3}+\frac{19}{10}\zeta_{4}+\frac{3}{8}\zeta_{5}\right)+\frac{d^{abc}_{F}d^{abc}_{F}}{N_{c}}\left(\frac{69}{40}+\frac{42}{5}\zeta_{3}-12\zeta_{5}\right)\,\,\,\bigg]~.\end{split}$ (7)
For finite fermion masses, both tensor structures are nonvanishing at Born level, $\eta_{\mu\nu}H^{\mu\nu}_{\mathcal{E}}=1+n_{\mu}n_{\nu}H^{\mu\nu}_{\mathcal{E}}=1+\frac{1}{2}(4m_{f}^{2}/q^{2})$. Near threshold $a_{\mathcal{E}}\to 0$ and the vector current behaves like a scalar operator. Compared with loop corrections of order $\alpha_{s}/\pi$, this effect from the mass is relevant for $q^{2}\lesssim 4m_{f}^{2}(\pi/\alpha_{s})$, and we shall see that it is sizable for $b$ quarks even away from threshold.
In QCD with $n_{f}=5$, all numerical coefficients of $\alpha_{s}^{k}$ in Eq. 7 are all positive. At $q^{2}=m_{Z}^{2}$ and using $\alpha_{s}(m_{Z}^{2})=0.118$, one gets $a_{\mathcal{E}}=-0.4437$, $a_{\mathcal{E}}=-0.4258$ and $a_{\mathcal{E}}=-0.42$ at NLO, NNLO and N3LO, respectively.
Even though there is no experimental measurement explicitly dedicated to the one-point energy correlator, we observe that the expressions in Eq. 6 directly relate $a_{\mathcal{E}}$ to existing measured observables through
$a_{\mathcal{E}}=-\frac{1}{2}+\frac{3}{2}\frac{\sigma_{L}}{\sigma_{\text{tot}}}~,$ (8)
where $\sigma_{L}$ is the longitudinal cross section in semi-inclusive $e^{+}e^{-}\to hX$. In this representation, the positivity constraints are simply $\sigma_{T,L}\geq 0$. The longitudinal ratio $\sigma_{L}/\sigma_{\rm tot}$ has been measured at LEP by both OPAL and DELPHI Collaborations in Akers and others (1995); Abreu and others (1999) at the $Z$ pole. The JADE Collaboration, at PETRA, also reported a measurement at $\sqrt{q^{2}}=36.6\,\text{GeV}$ Blumenstengel et al. (2001). The data at SLAC at $7.4\,\text{GeV}$ in Schwitters and others (1975); Feldman and Perl (1977) may also be interpreted in terms of $\sigma_{L}/\sigma_{\rm tot}$, although with statistical errors only, as done in Blumenstengel et al. (2001). The interpretation of these measurements in terms of $a_{\mathcal{E}}$ is summarized in Table 1.
The LEP measurements agree well with the N3LO results. The JADE value is compatible with the N3LO result of $a_{\mathcal{E}}=-0.396$ at $\sqrt{q^{2}}=36.6\,\text{GeV}$, obtained after including the $b$-quark mass effect. The SLAC measurement does not contain systematic uncertainties, but seems to validate the perturbative approach. Theoretical and parametric uncertainties will be discussed with the results in Fig. 1. It should be emphasized that, despite their equivalence, dedicated experimental measurements of $a_{\mathcal{E}}$ through the energy correlator should be simpler and cleaner than $\sigma_{L}/\sigma_{\rm tot}$ measurements.
Experiment $\sqrt{s}$ [GeV] $a_{\mathcal{E}}$ Ref.
DELPHI 91.2 $0.423\pm 0.011$ Abreu and others (1999)
OPAL 91.2 $0.415\pm 0.008$ Akers and others (1995)
JADE 36.6 $0.40\pm 0.02$ Blumenstengel et al. (2001)
SLAC 7.4 $0.35\pm 0.03$ Feldman and Perl (1977); Blumenstengel et al. (2001)
Table 1: Indirect determinations of $a_{\mathcal{E}}$ from $\sigma_{L}/\sigma_{\rm tot}$ measurements in $e^{+}e^{-}$ data at different center of mass energies.
[…]
The contributions to $a_{\mathcal{E}}$ from two and three pseudoscalar states, which dominate the cross section at low energy, are fixed by symmetries. Two-pseudoscalar states saturate the upper bound $a_{\mathcal{E}}=1$. This can be explicitly seen since the hadronic current is fixed to have the form $J^{\mu}=(p_{1}^{\mu}-p_{2}^{\mu})F^{V}_{P}(q^{2})$ where $p_{1,2}^{\mu}$ are the pseudoscalar momenta, and $F^{V}_{P}(q^{2})$ is a form factor that defines the meson’s charge via $F^{V}_{P}(0)=Q_{P}$. The matrix element leads to the relation $\eta_{\mu\nu}H^{\mu\nu}_{\mathcal{E}}=n_{\mu}n_{\nu}H^{\mu\nu}_{\mathcal{E}}$, saturating the upper bound of $a_{\mathcal{E}}$ independently of the form factor and pseudoscalar mass. It should be noted that $\alpha_{\text{em}}$ corrections to this extremal value, leading to final state radiation, can only decrease $a_{\mathcal{E}}$ due to unitarity.
For states with three pseudoscalars, the structure of the hadronic current is determined by parity considerations, given by $J_{\mu}=\epsilon_{\mu\nu\alpha\beta}p_{1}^{\nu}p_{2}^{\alpha}p_{3}^{\beta}\mathcal{F}(s_{1},s_{2},s_{3})$  Hoferichter et al. (2018), with $\mathcal{F}(s_{1},s_{2},s_{3})$ being some function of $s_{i}=(q-p_{i})^{2}$ fixed by the WZW term in the $s_{i}\to 0$ limit Adler et al. (1971); Aviv and Zee (1972). Independently of the form factor $\mathcal{F}(s_{i})$, the tensor structure implies that the projection of the correlator on $n^{\mu}n^{\nu}$ vanishes, $n_{\mu}n_{\nu}H^{\mu\nu}=0$, thus, saturating the lower bound $a_{\mathcal{E}}=-1/2$. The vector-pseudoscalar channels, like $\omega\pi$ and $\omega\eta$, are also fixed by symmetry and saturate the lower bound $a_{\mathcal{E}}=-1/2$, as it can be explicitly seen from the $\epsilon^{\mu\nu\lambda\delta}\partial_{\mu}\omega_{\nu}\partial_{\lambda}\rho^{i}_{\delta}\pi^{i}$ vertex Fujiwara et al. (1985). It is worth noticing how all currents leading to $a_{\mathcal{E}}=-1/2$ are induced by the WZW term.
[…]
The aforementioned channels dominate the cross section up to $\sqrt{q^{2}}\simeq 1.4\,\text{GeV}$, above which many other channels open. We take into account the contributions of these channels by extracting the exclusive contributions to the total hadronic $R$ ratio between $1.12\,\text{GeV}$ and $1.937\,\text{GeV}$ from the analysis in Keshavarzi et al. (2018). For these channels, notably for $KK\pi\pi$, $5\pi$ and $6\pi$, reliable predictions for $a_{\mathcal{E}}$ are currently unavailable. Therefore, they become the main source of uncertainty around $2\,\text{GeV}$. The impact of this uncertainty in the evolution of $a_{\mathcal{E}}$ will be detailed in the following discussion of Fig. 1.
Since the extremality of the two and three pseudoscalar channels crucially relies on the precise isolation of the initial state radiation, it provides a nontrivial consistency test of its treatment. This could have an impact on the assessment of their contribution to the hadronic vacuum polarization.
[…]
At large $q^{2}$, the perturbative QCD results, obtained using Eq. 4 with $\eta_{\mu\nu}H^{\mu\nu}_{\mathcal{E}}$ and $n_{\mu}n_{\nu}H^{\mu\nu}_{\mathcal{E}}$ are shown at NLO (yellow), at NNLO (teal) and at N3LO (red). In these, we consider the finite mass effects for the $b$- and $c$-quarks, only at Born level, while light quarks are treated as massless. The spread of lines for each result in Fig. 1 reflects their uncertainty, estimated by drawing the input parameters from a gaussian distribution around the current experimental ranges: $m_{b}=4.18\pm 0.03,\text{GeV}$, $m_{c}=1.27\pm 0.02,\text{GeV}$, and $\alpha_{s}(m_{Z}^{2})=0.118\pm 0.0016$ Navas and others (2024), and the renormalization scale $\mu$ uniformly over the range $\mu=\sqrt{q^{2}}\times 2^{\pm 1}$. The uncertainty of the NLO result is dominated by the scale variation. On the contrary, the N3LO uncertainty is dominated by the one associated with input parameters, mostly $\alpha_{s}(m_{Z}^{2})$. The grey vertical bands indicate presence of resonances in the charmonium and bottomonium region.
The four-loop calculation of the hadronic vacuum polarization Baikov et al. (2008) shows good agreement with nonresonant experimental data in the energy range $\sqrt{q^{2}}\in[1.8,3.7]\,\text{GeV}$ Davier et al. (2020). It is natural to expect a similar agreement for $a_{\mathcal{E}}$. This motivates the tentative extension of the perturbative results for $a_{\mathcal{E}}$ in Fig. 1 up to $\sqrt{q^{2}}\simeq 2\,\text{GeV}$. Given the difference between the mean value of the N3LO and the previous order, it is important to assess whether the $\alpha_{s}^{4}$ correction stabilizes. It would also be valuable to include $\Lambda_{QCD}/\sqrt{q^{2}}$ corrections.
The four data points in Fig. 1 represent the recast of the experimental measurements as summarized in Table 1. To our best knowledge, this is the first interpretation of data in terms of $a_{\mathcal{E}}$.
In the low energy region, the black curve from the pion threshold up to $1.12\,\text{GeV}$ shows the prediction from data-driven methods as obtained with the PHOKHARA Monte Carlo. The two-pion states saturate $a_{\mathcal{E}}=1$, while the $\omega$ and $\phi$ resonances give sharp features due to the sudden enhancement of the rate for tree-pion states, which saturate the $a_{\mathcal{E}}=-1/2$ boundary. Above the $\phi$ resonance, the four-pion states start to contribute, rapidly pushing down the value of $a_{\mathcal{E}}$.
Between $1.12\,\text{GeV}$ and $1.936\,\text{GeV}$ we use the rates as extracted from data in Keshavarzi et al. (2018) with their $a_{\mathcal{E}}$ predictions. For the channels that we do not have a reliable prediction for $a_{\mathcal{E}}$, like the $KK\pi\pi$, $5\pi$, $6\pi$, $\omega\eta\pi$ and others, we consider a conservative flat prior for $a_{\mathcal{E}}$ within the unitarity bounds. Drawing multiple samples from this prior leads to the spread of black lines in Fig. 1.
Around $2\,\text{GeV}$, both high and low energy approaches are at the limits of their validity. Nonetheless, they give compatible predictions for $a_{\mathcal{E}}$, giving a coherent global picture for its evolution in QCD.
[…]
The one-point energy correlator, by virtue of its theoretical and experimental simplicity, is a clean and powerful probe of QCD across scales, and a new benchmark for testing the emergence of hadronic degrees of freedom.
Acknowledgments—  We thank Gauthier Durieux, Heejoo Kim, Lorenzo Ricci and Francesco Riva for discussions, suggestions and comments on the draft. MS was supported by National Research Foundation of Korea under Grant Number RS-2024-00450835.